OpenStudy (howard-wolowitz):

Find a general formula for the nth term of the geometric series with:

OpenStudy (howard-wolowitz):

\[a _{9}=1/7.......r=1/7\]

OpenStudy (howard-wolowitz):

@SolomonZelman

OpenStudy (solomonzelman):

is that \(\Large\color{black}{ \displaystyle a_{9}}\) ?

OpenStudy (howard-wolowitz):

yes

OpenStudy (solomonzelman):

so you have a common ratio 1/7 and 1st term of 1/7. now, for any geom. sequence, you know: \(\Large\color{black}{ \displaystyle a_n=a_{k}\times(r)^{n-k}}\) WHERE: (1) k and n are natural numbers can be denoted as \(\Large\color{black}{ \displaystyle \left\{ k,~n\right\}\in{\bf N}}\) (2) \(\Large\color{black}{ \displaystyle n>k}\) n is a bigger natural number than k.

OpenStudy (solomonzelman):

\(\Large\color{black}{ \displaystyle a_9=a_{1}\times\left(\frac{1}{7}\right)^{9-1}}\)

OpenStudy (howard-wolowitz):

so \[a _{n}=7^{6,7,8,9} (1/7) ^{n-1}\]

OpenStudy (solomonzelman):

\(\Large\color{black}{ \displaystyle\frac{1}{7}=a_{1}\times\left(\frac{1}{7}\right)^{9-1}}\)

OpenStudy (solomonzelman):

so, this way you can find the 1st term

OpenStudy (solomonzelman):

then write, \(\Large\color{black}{ \displaystyle a_n=a_1\times (r)^{n-1}}\) formula, BUT FILL IN YOUR VALUES

OpenStudy (solomonzelman):

or do you want a \(\Large\color{black}{ \displaystyle \sum}\) notation ?

OpenStudy (howard-wolowitz):

well the only i gotta do now is find what the exponent is

OpenStudy (howard-wolowitz):

also sorryu no i dont

OpenStudy (solomonzelman):

find the exponent ?

OpenStudy (howard-wolowitz):

so after i fill in my values are you thinking its 6 or 7

OpenStudy (howard-wolowitz):

well see what i wrote above yours ^

OpenStudy (howard-wolowitz):

it has to be one of those exonents

OpenStudy (howard-wolowitz):

and I think its 7

OpenStudy (solomonzelman):

have you found the first term?

OpenStudy (howard-wolowitz):

dont have to its just \[a\]

OpenStudy (solomonzelman):

\(\Large\color{black}{ \displaystyle \frac{1}{7}=(a_1) \times \left( \frac{1}{7} \right)^{8} }\) \(\Large\color{black}{ \displaystyle 7^{1}=(a_1) \times \left( 7^{-1} \right)^{8} }\) \(\Large\color{black}{ \displaystyle 7^{1}=(a_1) \times \left( 7^{-8} \right)}\) \(\Large\color{black}{ \displaystyle 7^{-1} \div 7^{-8}=a_1 }\) \(\Large\color{black}{ \displaystyle 7^{-1 -(-8)}=a_1 }\) \(\Large\color{black}{ \displaystyle 7^{7}=a_1 }\)

OpenStudy (solomonzelman):

this si what I got

OpenStudy (solomonzelman):

is*

OpenStudy (solomonzelman):

and your common ratio is \(\Large\color{black}{ \displaystyle 7^{-1} }\)

OpenStudy (solomonzelman):

just a?

OpenStudy (howard-wolowitz):

so i was right? I thought it was 7 too

OpenStudy (howard-wolowitz):

\[a _{n}\]

OpenStudy (solomonzelman):

don't know what you meant by your last comment. but your terms are so, as you said your terms will just be \(\Large\color{black}{ \displaystyle 7^{7},~~~~ 7^{6},~~~~7^{5},~~~~ 7^{4},~~~~}\) etc... \(\Large\color{black}{ \displaystyle \Downarrow,~~~~~ \Downarrow,~~~~~\Downarrow,~~~~~ \Downarrow,~~~~}\) etc... \(\Large\color{black}{ \displaystyle a_1,~~~~ a_2~~~~a_3,~~~~ a_4,~~~~}\) etc...

OpenStudy (howard-wolowitz):

but we agree its \[7^{7}\]

OpenStudy (solomonzelman):

So, your formula \(\Large\color{brown}{ a_{\rm n}= a_{\rm 1} \times ({\rm r})^{{\rm n}-1} }\) can be written as \(\Large\color{blue }{ a_{\rm n}= \left(7^7 \right) \times ({\rm 7^{-1}})^{~{\rm n}-1} }\)

OpenStudy (solomonzelman):

or you can write 7^(-1) as 1/7, it is all same./

OpenStudy (solomonzelman):

the first term, yes

OpenStudy (solomonzelman):

or, if you want to add all the terms, then write \(\Large\color{blueviolet}{ \displaystyle \sum_{ n=1 }^{ \infty } 7^{(8-n)}}\)

OpenStudy (solomonzelman):

my connection is very bad, sorry for delays

OpenStudy (howard-wolowitz):

same here dude